August 2015

Report Date
August 2015

TEL Programme Report compiled by Colin Watt, (IS Apps Programme Manager) for Mark Wetton and Anne-Marie Scott

 

Section One: Programme Commentary, RAG and Resources Update

Section Two: Projects Not Started

Section Three: Projects In Progress

Section Four: Closed or Withdrawn Projects 

 

Section One: Programme Commentary, RAG and Resources Update

 

Overall Programme RAG: GREEN

The Programme starts the year in a strong position, albeit with a number of 14/15 projects closures pushed into 15/16.

TEL017 VLE Student Analytics - now ready to close, approval expected early September.

TEL018 QMP Upgrade - DSOR and closure delayed due to prolonged deployment and resolution of technical issues, closure now expected in September, no further delays anticipated.

TEL021 TELS Management Information - 

TEL022 Moodle Upgrade - now ready to close, approval expected early September.

TEL023 Next Generation Personal Response System Implementation -  now ready to close, approval expected early September.

TEL024 Migration of MVM UG to PebblePad - due for completion September.

TEL025 Moodle Data Feeds - Planning completion due mid September; scope and cost is subject to review by Architecture group of the proposal.

 

Forward Look and RAG

 

Resource Summary

 

 

 

Section Two: Projects Not Started

 

Project InfoProject Manager's Commentary
AP56-010 BlackBoard Collaborate MyEd Channel Refresh

Scheduled start: August 2015; kickoff meeting has taken place online; requires PM to be assigned.

AP56-012 Coldfusion Upgrade of Learn Services MyEd ChannelScheduled start: January 2016
AP56-013 Annual Moodle Upgrade

Scheduled start: January 2016

AP56-079 VLE Student AnalyticsScheduled start: January 2016
AP56-080 Online Assessment Evaluations (years 2 and 3)Scheduled start: tbd
AP56-035 Learn and Student Record Grade ExchangeScheduled start: October 2015; PRICCL required to transfer 10 days effort from the TEL programme to the SSP Programme to cover R&D under their existing project. 

 

 

 

 

Section Three: Projects in Progress

 

Project InfoProject Manager's Commentary

TEL015 (AP45-039)

Online Assessment/Feedback Evaluations – FACS

Milestones:

 

01-Oct-2015End of suspension

 

RAG Status

 

SUSPENDED

 

Report by Sue Woodger for April 2015

 

TEL015 suspended for 6 months until 1 October 2015.

 

Next Steps

End of Suspension 1 Oct 15

 

TEL017 (AP45-038)

VLE Student Analytics

Milestones

 
31-Aug-2015Closure

RAG Status

GREEN

Report by Morna Findlay for August 2015  

Achievements in the Last Period

  • UAT completed
  • Performance testing completed with successful results
  • Acceptance review held and approved
  • Deployment to LIVE completed
  • SDD and TAD updated and approved
  • Handover to production completed
  • Deployment sign off review held and approved
  • Closure report approved by email
  • Project to be submitted for closure approval WIS 04/09/15

TEL018 (AP45-037)

QMP Upgrade

Milestones:

  
11-Sep-2015Closure

 

RAG Status

GREEN

Report by Colin Forrest for August 2015  

Achievements in the Last Period

  • The project deployment has been signed off and we are ready to close the project.

Issues

  • The upgrade encountered many issues during the progress of the project.  However the team are pleased to have the supported v5.7 in production.  It is a better position for the University to be in and we can now work through our remaining issues with the supplier.
  • We have gone 6 days over our original 100 day budget.

Next steps

  • The closure report will be completed and the project closed.

TEL021 (AP45-039)

TELS Management Information

Discretionary  

 

Milestones 

 

11-Sep-2015Acceptance sign-off [Moodle]
30-Sep-2015Delivery [Moodle]
08-Oct-2015Deployment sign-off [All]
15-Oct-2015Closure

 

RAG Status

AMBER

Report by David Watters for August 2015

 

 

Achievements in Last Period

  • Live deployment of Learn, PebblePad & QMP elements
  • Moodle parts of project ready for user testing

Issues

  • Project effort has increased; still to be agreed with Project Sponsor
  • Learn elements not available in Test because of permissions problem

Next Steps

  • UAT of Moodle elements
  • Deploy Moodle elements into Live
  • Eradicate minor issues in Live for non-Moodle elements

 

RAG status is AMBER because the additional effort has not yet been signed off as agreed.

TEL022 (AP45-058)

Moodle Upgrade

Compliance  

 

Milestones   

14-Aug-2015DSOR (Delivery Sign Off)
28-Aug-2015Closure

 

RAG Status

GREEN

Report by Colin Watt for August 2015

Achievements in Last Period

  • DSOR completed and signed off.
  • Closure survey and report completed for submission w/c 07/09/15.

 

Next Steps

Complete Closure.

 

Issues

None.

TEL023 (AP34-049)

Next Generation Voting System Implementation

Compliance  

 

Milestones   

09-Sep-2015Project Closure

 

RAG Status

GREEN

Report by Maurice Franceschi for August 2015

 

The service had a soft launch on 2nd July and  once contract signed and some fixes made - formally launched on 28th July.

Achievements in Last Period

  • SSO for applications from TopHat pending in July needs checked - done
  • follow up on the rooms identified as at risk for insufficient capacity - done
  • follow up on NHS staff using Top Hat. - done - no further action here. Paul Clark informed.
  • decommission clickers system - done. SOftware removed from managed desktop.

Issues

  • none
  • no amber or red risks

Next Steps

closure

 

TEL024

TEL024 Migration of MVM UG to PebblePad

Discretionary

Milestones

 

  
11-Sep-2015Year 2 ready for submissions
11-Sep-2015Year 1 ready for submissions
25-Sep-2015Closure

 

 

RAG Status

GREEN

 

Report by Sue Woodger for August 2015

 

Achievements in Last Period

  • Year 3 ready for submissions
  • Clinical skills log and the ‘Big Five’ Clinical skills submissions are now live. Guidance for these items has also been created in EEMeC
  • Year 4 MBChB Workspace complete - assignments set up, feedback templates attached and ready to go
  • Update for Years 4 and 5 students went out 18 August clarifying points following their first submission (there was a small number of students who had not read the guidance on submitting and had issues – this update confirms process for all students for portfolio and skills )
  • Training for year 1 co-ordinators completed 19 August
  • Training for year 3 co-ordinators completed 20 August
  • Training for year 2 co-ordinators completed 21 August
  • The NHS Lothian access issue has now been resolved. This was due to a NHS Lothian security issue which prevented downloads in Internet Explorer. However all machines seemed to have Firefox installed so can be managed by communication rather than requiring any additional software etc. MVM has set up a staff FAQ for markers on EEMeC

Issues

  • NHS Borders issue still ongoing - currently looks like we don’t have access to the URL / website at all in NHS Borders. Lydia liaising directly with NHS Borders – they are looking into getting URLs unblocked.

Next Steps

  • Year plans to be signed off this week
  • Year 2 ready for submissions by 11 September
  • Year 1 ready for submissions by 11 September
  • Documenting the processes for handover
  • Closure report

 

The project status is GREEN

 

TEL025

TEL025 Moodle Data Feeds

Discretionary

Milestones

 

14-Aug-2015Project Brief completed
02-Oct-2015Business Analysis complete (TBC)
30-Oct-2015Design Complete (TBC)
11-Dec-2015Build Complete (TBC)
18-Dec-2015Integration Complete (TBC)
12-Feb-2016Acceptance complete (TBC)
18-Feb-2016Deployment complete (TBC)
25-Feb-2016Deployment Review complete (TBC)
04-Mar-2016Closure (TBC)

RAG Status

GREEN

Report by Karen Stirling August 2015

Achievements in Last Period

  • Project stakeholders identified
  • High level project plan created
  • Project brief drafted and issued out for review by project team and project sponsor
  • Project brief review meeting held to estimate resource

 

Issues

  • Budget for project has been limited to 50 days following annual planning as advised by Project Sponsor
  • At project brief review it was agreed that this project should be reviewed by the Applications Architectural Group as there are a number of projects looking to complete similar deliverables e.g. SAC034 - Assessment and Progression in EUCLID and HSS are looking for interfaces to / from EUCLID.
  • Project team agree at a high level that the project estimate will be in the region of 70 days based on a similar project for the Library IDM Feed. This information has been passed to the Programme Manager and Project Sponsor to consider where additional funding can be acquired for the project.
  • Further clarity on business requirements needed to allow project brief to be submitted, Service Manager on leave however this information is now available to include in brief

 

 Next Steps

  • Planning and Resource Estimates
  • Project Brief approval

 

The project status is GREEN

 

Section Four: Closed or Withdrawn Projects

 

 

Project InfoProject Managers Commentary